Grade 9Algebra

root (root ( 7 + root 48)) – (root (7 – root 48) )) =

ra
11 Years agoGrade 9

2 Answers

Akshay
11 Years ago
Let , root (root ( 7 + root 48)) – (root (7 – root 48) )) = x,
square, root ( 7 + root 48) – root (7 – root 48) = x2,
square again, 7+rt48 + 7-rt48 -2rt(49-48) = x4 ,
12=x4,
ATB
Lakshay
9 Years ago
7+root 48 = 7 + root 2×2×2×2×2×3 so 7+root 48= 7+ 4 root 3Using a+b hole square and middle term spliting answer will beRt3square +rt2square + 2×2×rt 3= 2+rt3 hole square